OpenStudy (anonymous):

EVALUATE LIM 3XSIN(X) + COS (X) / X^2 AS X GOES TO INFINITY

OpenStudy (amistre64):

The equation is messy, it is hard to tell exactly what it is trying to say.

OpenStudy (anonymous):

\[3x \sin (x) + \cos (x) / x ^{2}\]

OpenStudy (anonymous):

the answer is 0 but how to support my work ??

OpenStudy (amistre64):

hmm..... try dividing everything by x^2 the bottom goes to 1 so we aint got to worry about it. 3xsin(x)/x^2 + cos(x)/x^2 3sin(x)/x + cos(x)/x^2 Anything with an x left in the bottom goes to 0 as x-->infinity. (0+0)/1 = 0. Thats the best I got for that one :)

OpenStudy (anonymous):

thank you